OpenStudy (anonymous):

Solve 24 - 3x = -27.

jigglypuff314 (jigglypuff314):

subtract both sides by 24 divide both sides by negative 3

OpenStudy (anonymous):

x=20 with a remander of 3

jigglypuff314 (jigglypuff314):

that would not be correct...

jigglypuff314 (jigglypuff314):

24 - 3x = -27 subtract both sides by 24 -24 -24 -3x = -51 divide both sides by negative 3 /-3 /-3 x = 51/3 = (simplify)

OpenStudy (anonymous):

woops i wrote 6 instead of 5.. thats why it was wrong

OpenStudy (anonymous):

Thank you